# On the Distribution of Points of Valuation $1$ for a Polynomial in Two Variables

### v1

@@ 1 Introduction
One essential result in $p$-adic valuation is Legendre’s formula[7], which calculates the $p$-adic valuation of $n!$, For a prime $p$, the formula is given by:
$\nu_{p}(n!)=\sum_{k=1}^{\infty}\left\lfloor\frac{n}{p^{k}}\right\rfloor$
This sums the number of multiples of $p,p^{2},p^{3},\dots$ less than or equal to $n$, capturing how many times $p$ divides $n!$. Introduced by Adrien-Marie Legendre in the 19th century, this formula has important applications in combinatorics, particularly in understanding binomial coefficients and their divisibility by primes.
@@
for a given sequence $x=\{x_{n}\}$ of integers.
The $2$-adic valuation of $\{n^{2}-a\}$ is studied in [3]. It is shown that $n^{2}-a,a\in\mathbb{Z}$ has a simple closed form[2, 4] when $a\not\equiv 4,7\mod 8$. For these two remaining cases the valuation is quite complicated. It is studied by the notion of the valuation tree. In [9], the set
@@
Denote $C_{p}=\{(x,y)\in\mathbb{Z}^{2}:\nu_{p}(f(x,y))>0\textit{ and }0\leq x,y<p\}$ be the points of the curve given by the equation $f(x,y)=0$ over $\mathbb{F}_{p}$ and $m=|C_{p}|$ be the number of points in $C_{p}$.
@@
These seem to encode some elements of randomness which we try to quantify further by studying the properties of indicator random variables $Y_{x,y}$ defined in Section 3. To summarize, these variables are identically distributed Bernoulli random variables with probability $1/p$ and are pairwise “nearly” independent.
In a different direction, we study the variation in total number of points with fixed valuation in a random square. More precisely, define the random variable $X:\mathbb{F}_{p}\times\mathbb{F}_{p}\rightarrow\mathbb{Z}$ where $X(k,l)$ is the total number of points $(x,y)\in\{k,k+1,\ldots,k+p-1\}\times\{l,l+1,\ldots,l+p-1\}$ such that $\nu_{p}(f(x,y))>1$. In other words, $X$ represents the number of points in the random $p\times p$ square where the valuation of $f$ is greater than $1$. Observing the distribution of $X$ for various primes $p$, we get the following Figure 2.
@@
Before stating Conjecture 1, we would need some notations. For any $k$-tuple $((r_{1},s_{1}),$ $(r_{2},s_{2}),\dots,(r_{k},s_{k}))$ of points on the curve $C_{p}$, denote the associated matrix $B_{k}$ to be
$B_{k}=\begin{bmatrix}f_{x}(r_{1},s_{1})&f_{y}(r_{1},s_{1})&f(r_{1},s_{1})/p\\ f_{x}(r_{2},s_{2})&f_{y}(r_{2},s_{2})&f(r_{2},s_{2})/p\\ \vdots&\vdots&\vdots\\ f_{x}(r_{k},s_{k})&f_{y}(r_{k},s_{k})&f(r_{k},s_{k})/p\end{bmatrix}.$ (1)
Let $m_{k,1}$ denote the number of element-wise distinct $k$-tuples of points on $C_{p}$ whose associated mattrix $B_{k}$ has rank $1$. Let $m_{k,2}$ denote the number of element-wise distinct $k$-tuples for which $B_{k}$ has rank $2$ and the first $2$ columns of $B_{k}$ are linearly independent.
@@ 3 Indicator random variables for valuation greater than $1$
Recall from Section 1, that $m_{k,2}$ denotes the number of element-wise distinct $k$-tuples of points on $C_{p}$ whose associated matrix $B_{k}$ given by (1) has rank $2$ with the first $2$ columns of $B_{k}$ being linearly independent and $m_{k,1}$ denote the number of element-wise distinct $k$-tuples where $B_{k}$ has rank $1$.
@@
$B_{k}\begin{pmatrix}i\\ j\\ 1\end{pmatrix}\equiv\begin{pmatrix}0\\ 0\\ \vdots\\ 0\end{pmatrix}\pmod{p}.$ (15)
In case the rank of $B_{k}$ is $2$ and the first $2$ columns are linearly independent, there is a unique solution of (15) and hence $n=1$. In case the rank of $B_{k}$ is $1$, there are $n=p$ distinct solutions of (15). In the other cases, either rank of $B_{k}$ is $3$ or the rank of $B_{k}$ is $2$ and the first $2$ columns are linearly dependent, it can be seen that there are no solutions of (15).

### v2

@@ 1 Introduction
One essential result in $p$-adic valuation is Legendre’s formula[7], which calculates the $p$-adic valuation of $n!$. For a prime $p$, the formula is given by:
$\nu_{p}(n!)=\sum_{k=1}^{\infty}\left\lfloor\frac{n}{p^{k}}\right\rfloor\cdot$
This sums the number of multiples of $p,p^{2},p^{3},\dots$ less than or equal to $n$, capturing how many times $p$ divides $n!$. This formula has important applications in combinatorics, particularly in understanding binomial coefficients and their divisibility by primes.
@@
for a given sequence $x=\{x_{n}\}$ of integers. The $2$-adic valuation of $\{n^{2}-a\}$ is studied in [3]. It is shown that $n^{2}-a,a\in\mathbb{Z}$ has a simple closed form[2, 4] when $a\not\equiv 4,7\mod 8$. For these two remaining cases the valuation is quite complicated. It is studied by the notion of the valuation tree. In work by the second author [9], the set
@@
For a fixed polynomial $f(x,y)\in\mathbb{Z}[x,y]$, let $C_{p}=\{(x,y)\in\mathbb{Z}^{2}:\nu_{p}(f(x,y))>0\textit{ and }0\leq x,y<p\}$ be the points of the curve given by the equation $f(x,y)=0$ over $\mathbb{F}_{p}$ and $m=|C_{p}|$ be the number of points in $C_{p}$.
@@
These seem to encode some elements of randomness which we try to quantify further by studying the properties of indicator random variables $Y_{x,y}$ defined in Section 3. To summarize, these variables are identically distributed Bernoulli random variables with probability $1/p$ and are pairwise independent most of the time.
In a different direction, we study the variation in total number of points with fixed valuation in a random square. More precisely, define the random variable $X:\mathbb{F}_{p}\times\mathbb{F}_{p}\rightarrow\mathbb{Z}$ where $X(k,l)$ is the total number of points $(x,y)\in\{kp,kp+1,\ldots,kp+p-1\}\times\{lp,lp+1,\ldots,lp+p-1\}$ such that $\nu_{p}(f(x,y))>1$. In other words, $X$ represents the number of points in the random $p\times p$ square where the valuation of $f$ is greater than $1$. Observing the distribution of $X$ for various primes $p$, we get the following Figure 2.
@@
Before stating Conjecture 1, we would need some notations. For any $k$-tuple $\mathbf{P}=((r_{1},s_{1}),(r_{2},s_{2}),\dots,(r_{k},s_{k}))$ of points on the curve $C_{p}$, denote the associated matrix $B_{k}(\mathbf{P})$ to be
$B_{k}(\mathbf{P})=\begin{bmatrix}f_{x}(r_{1},s_{1})&f_{y}(r_{1},s_{1})&f(r_{1},s_{1})/p\\ f_{x}(r_{2},s_{2})&f_{y}(r_{2},s_{2})&f(r_{2},s_{2})/p\\ \vdots&\vdots&\vdots\\ f_{x}(r_{k},s_{k})&f_{y}(r_{k},s_{k})&f(r_{k},s_{k})/p\end{bmatrix}.$ (1)
Let $m_{k,1}$ denote the number of element-wise distinct $k$-tuples of points on $C_{p}$ whose associated matrix $B_{k}(\mathbf{P})$ has rank $1$. Let $m_{k,2}$ denote the number of element-wise distinct $k$-tuples for which $B_{k}(\mathbf{P})$ has rank $2$ and the first $2$ columns of $B_{k}(\mathbf{P})$ are linearly independent.
@@ 3 Indicator random variables for valuation greater than $1$
Recall from Section 1, that $m_{k,2}$ denotes the number of element-wise distinct $k$-tuples of points on $C_{p}$ whose associated matrix $B_{k}(\mathbf{P})$ given by (1) has rank $2$ with the first $2$ columns of $B_{k}(\mathbf{P})$ being linearly independent and $m_{k,1}$ denote the number of element-wise distinct $k$-tuples where $B_{k}(\mathbf{P})$ has rank $1$.
@@
$B_{k}(\mathbf{P})\begin{pmatrix}i\\ j\\ 1\end{pmatrix}\equiv\begin{pmatrix}0\\ 0\\ \vdots\\ 0\end{pmatrix}\pmod{p}.$ (15)
In case the rank of $B_{k}(\mathbf{P})$ is $2$ and the first $2$ columns are linearly independent, there is a unique solution of (15) and hence $n=1$. In case the rank of $B_{k}(\mathbf{P})$ is $1$, there are $n=p$ distinct solutions of (15). In the other cases, either rank of $B_{k}(\mathbf{P})$ is $3$ or the rank of $B_{k}(\mathbf{P})$ is $2$ and the first $2$ columns are linearly dependent, it can be seen that there are no solutions of (15).